OpenStudy (anonymous):

Use the Pythagorean theorem to prove the Pythagorean identity: (given: a squared + b squared = c squared, sin of data = a/c, cos of data = b/c

OpenStudy (dan815):

u want prof of pythagorean thrm or sin theta = a/c and cos theta = b/c

OpenStudy (anonymous):

Both

OpenStudy (dan815):



OpenStudy (dan815):

start with this

OpenStudy (anonymous):

Ok

OpenStudy (dan815):

lots of nice proofs there